Product A consists of 20% phosphoric acid, 12% sodium polyacrylate,
15% potassium pyrophosphate and 53% water (all wt%). There are 1,500
LBS of product A in stock.
It has been decided that Product A would be eliminated from the
company's product line and replaced with Product B.
Product B consists of 3% zinc sulfate, 25% phosphoric acid, 8% sodium
polyacrylate and 19% potassium pyrophosphate (all wt%).
We want to make 4000 LBS of product B, and we want to use Product A as
a base since it already contains most of the ingredients. How many LBS
of the following ingredients do we need to add to 1500 LBS of Product A
in order to make 4000 LBS of product B?
a. Zinc Sulfate
b. Phosphoric Acid
c. Sodium Polyacrylate
d. Potassiu
1
Expert's answer
2020-07-30T07:06:43-0400

a) m(ZnSO4) = 4000*0.03= 120 LBS

b) m(H3PO4) = 4000*0.25-1500*0.2=700 LBS

c) m (sodium polyacrylate) = 4000*0.08-1500*0.12=140 LBS

d) m(potassium pyrophosphate) = 4000*0.19-1500*0.15=535 LBS.